OpenStudy (anonymous):

Quadratic formula on the docket! Any math lovers able to make sense of this problem for me? 6x^2-9x+1=0 I'm supposed to solve using the quadratic formula which is messing me up because I know how solve by completing the square.

OpenStudy (anonymous):

(but also rusty on the completing the square process)

OpenStudy (mokeira):

\[6x ^{2}-9x +1=0\] \[3x(2x-3) +1=0\] \[3x(2x-3)=-1\] \[2x-3=\frac{ -1 }{ 3x }\] \[\frac{ 2x }{ 1 }+ \frac{ 1 }{ 3x }=3\] \[2\frac{ 1 }{ 3 }x=3 \] \[x=3\div \frac{ 7 }{ 3 }\] \[x=\frac{ 9 }{ 7 } =1\frac{ 2 }{ 7 }\] I hope I am correct

OpenStudy (dumbcow):

@Mokeira \[\frac{2x}{1} + \frac{1}{3x} \ne 2 \frac{1}{3} x\]

OpenStudy (dumbcow):

@LaylaJ the quadratic formula and completing the square are same thing, the formula is derived using completing the square

OpenStudy (dumbcow):

quadratic formula is easy because its already solved for you, just plug in the numbers

OpenStudy (anonymous):

Ok, so I already know the basics of the qf since I know how to complete the square?

OpenStudy (dumbcow):

yes

OpenStudy (mokeira):

@dumbcow... true. i messed up there

OpenStudy (anonymous):

Guess I'll go back to my math book with renewed confidence. Thanks!

OpenStudy (dumbcow):

\[x = \frac{-b \pm \sqrt{b^2 -4ac}}{2a}\] a = 6 b = -9 c = 1

OpenStudy (anonymous):

lightbulb! :)

OpenStudy (anonymous):

Thank you for teaching me!